# Influences of non-standard interactions on PeV neutrino events with and without a $L_{\alpha}-L_{\beta}$ symmetry

### v1

@@ IV.1 Influences on the flavor ratio of high energy astrophysical neutrinos
Figure 4: Ternary plot of the flavor ratio at Earth in the $L_{\alpha}-L_{\beta}$ symmetric models. The fixed neutrino energy is 1PeV, Neutrino oscillation parameters are taken in the $3\sigma$ allowed range of $NuFIT$ 5.0 for the global fit data(NO). The left panel: $\mu^{\pm}$ damping source. Right panel: $\pi^{\pm}$ decay source. The orange dashed lines encompasses the 2020 $3\sigma$ C.R. with the $\pi$ decay source based on IceCubeAartsen et al. (2021). The purple dashed line covers the 2040 $3\sigma$ C.R. based on IceCube and IceCube-Gen2Aartsen et al. (2021). The black dash lines denotes the $3\sigma$ C.R. boundary based on the TeV-PeV neutrino telescopes available in 2040Song et al. (2021).
@@
For $\mu^{\pm}$ damping sources, the predictions of both the $L_{\alpha}-L_{\beta}$ symmetric and breaking models, deviate noticeably from the result of the standard model. The ratio in this case is sensitive to the neutrino oscillation and new physics parameters, and spans a noticeable area. In contrast, for $\pi^{\pm}$ decay sources, the flavor ratios of PeV neutrinos are almost focused on special points.
@@ IV.3 Influences on the events of PeV neutrinos
Here the spectral index $\alpha$ is related to the neutrino acceleration mechanism. We let $\alpha$ be equal to 2 Anchordoqui et al. (2014a, b). Using the best fit data of the netrino oscillation parameters (NO), the potential parameter $V_{\alpha\beta}$ and the diffuse flux $\phi_{\nu+\overline{\nu}}$, the ratio of the Glashow resonance event is calculated near the resonance energy (see Tab.4-5). For comparison purposes, the results obtained from the standard flavor conversion matrix are also listed in the tables.
@@
We can see that the Glashow resonant event ratio is obviously decreased by the NSI for neutrinos from $\mu^{\pm}$ damping source. At the best fit value of $V_{e\tau}$, the resonant event ratio in the symmetric model is around 30 times smaller than the value in SM scheme, while in the symmetry breaking model the ratio is near 5 times smaller than the standard result. Furthermore, we note that the ratio is sensitive to the parameter $V_{e\tau}$. Under the 95% upper limit, the ratio can be damped to $10^{-3}$ times the magnitude of the SM result. In contrast, the ratio is moderately reduced by the NSI for neutrinos from $\pi^{\pm}$ decay source. These observations are consistent with those read from Fig.6-7.

### v2

@@ IV.1 Influences on the flavor ratio of high energy astrophysical neutrinos
Figure 4: Ternary plot of the flavor ratio at Earth in the $L_{\alpha}-L_{\beta}$ symmetric models. The fixed neutrino energy is 1PeV, Neutrino oscillation parameters are taken in the $3\sigma$ allowed range of $NuFIT$ 5.0 for the global fit data(NO). The left panel: $\mu^{\pm}$ damping source. Right panel: $\pi^{\pm}$ decay source. The orange dashed lines encompasses the 2020 $3\sigma$ C.R. with the $\pi$ decay source based on IceCubeAartsen et al. (2021). The purple dashed line covers the 2040 $3\sigma$ C.R. based on IceCube and IceCube-Gen2Aartsen et al. (2021). The black dash lines denotes the $3\sigma$ C.R. boundary based on the TeV-PeV neutrino telescopes available in 2040Song et al. (2021). The black heart symbols represent the best-fit points.
@@
For $\mu^{\pm}$ damping sources, the predictions of both the $L_{\alpha}-L_{\beta}$ symmetric and breaking models, deviate noticeably from the result of the standard model. In contrast, for the models under the $\pi^{\pm}$ decay sources, the flavor ratio of PeV neutrinos is almost concentrated at special points. In particular, in the symmetric case (See right panel of Fig.4)) the regions of $L_{e}-L_{\tau}$ and $L_{\mu}-L_{\tau}$ case overlap each other.
Let us give a simple comment on the difference. When the potential parameter $V_{\alpha\beta}$ reaches a threshold, it would dominate the Hamilton and the flavor transition matrix could be fixed on a decoupling pattern ((see Eqs.17- 18) in the following section). Correspondingly, the flavor ratio at Earth could be given on a special point. Based on the ranges of parameter values in Tab.2-3, we can see that the parameters with the $\pi^{\pm}$ decay source approximate the flavor decoupling thresholds both in the $L_{\alpha}-L_{\beta}$ symmetric and breaking models. Hence the concentration of flavor ratio region appears.
@@ IV.3 Influences on the events of PeV neutrinos
Here $r_{e}$ is fraction of $\overline{\nu}_{e}$ at Earth, the spectral index $\alpha$ is related to the neutrino acceleration mechanism. We let $\alpha$ be equal to 2 Anchordoqui et al. (2014a, b). Using the best fit data of the netrino oscillation parameters (NO), the potential parameter $V_{\alpha\beta}$ and the diffuse flux $\phi_{\nu+\overline{\nu}}$, the ratio of the Glashow resonance event is calculated near the resonance energy (see Tab.4-5). For comparison purposes, the results obtained from the standard flavor conversion matrix are also listed in the tables.
@@
We can see that the Glashow resonant event ratio is obviously decreased by the NSI for neutrinos from $\mu^{\pm}$ damping source. At the best fit value of $V_{e\tau}$, the resonant event ratio in the symmetric model is around 30 times smaller than the value in SM scheme, while in the symmetry breaking model the ratio is near 5 times smaller than the standard result. Similar to the discussions on the flavor ratio in the previous section, the difference results from the Hamiltonian under the two models, which affects the flavor transition matrix. Using the best fit value of the $V_{e\tau}$ parameter, the flavor conversion matrices of the two models are obtained as follows:
$\overline{P}^{sm}(V_{e\tau})\simeq\left(\begin{array}[]{ccc}0.98{}&~{}0.0058{}&~{}0.0048{}\\ 0.0058{}&~{}0.65{}&~{}0.34{}\\ 0.0048{}&~{}0.34{}&~{}0.65\end{array}\right),~{}~{}\overline{P}^{bm}(V_{e\tau})\simeq\left(\begin{array}[]{ccc}0.48{}&~{}0.034{}&~{}0.48{}\\ 0.034{}&~{}0.94{}&~{}0.016{}\\ 0.48{}&~{}0.016{}&~{}0.49\end{array}\right),$ (20)
where $\overline{P}^{sm}(V_{e\tau})$ and $\overline{P}^{bm}(V_{e\tau})$ represent the matrix under the symmetric and symmetry breaking model respectively. We can see that the former matrix with the $\mu^{\pm}$ damping source can reduce the fraction $r_{e}$ to the order $10^{-3}$, while the latter gives $r_{e}$ of order $10^{-2}$. The damping of the Glashow resonance event ratio is proportional to the damping of $r_{e}$. Thus, the difference takes place. Furthermore, we note that the ratio is sensitive to the parameter $V_{e\tau}$. Under the 95% upper limit, the ratio can be damped to $10^{-3}$ times the magnitude of the SM result. In contrast, the ratio is moderately reduced by the NSI for neutrinos from $\pi^{\pm}$ decay source. These observations are consistent with those read from Fig.6-7.